# rem $\diamond$

### v1

@@ Abstract
The Fundamental Theorem of Statistical Learning states that a hypothesis space is PAC learnable if and only if its VC dimension is finite. For the agnostic model of PAC learning, the literature so far presents proofs of this theorem that often tacitly impose several measurability assumptions on the involved sets and functions. We scrutinize these proofs from a measure-theoretic perspective in order to extract the assumptions needed for a rigorous argument. This leads to a sound statement as well as a detailed and self-contained proof of the Fundamental Theorem of Statistical Learning in the agnostic setting, showcasing the minimal measurability requirements needed. We then discuss applications in Model Theory, considering NIP and o-minimal structures. Our main theorem presents sufficient conditions for the PAC learnability of hypothesis spaces defined over o-minimal expansions of the reals.
Math Subject Classification (2024): Primary 68T05, 03C64; Secondary 28A05, 28A20, 03C98, 68T27, 12J15. Key Words: PAC Learning, VC Dimension, Uniform Convergence, Measurability, Borel Sets, NIP, o-Minimality.
@@ 1 Introduction
In 1984, Valiant [30] introduced a formal model of learning based on randomly generated examples, which is referred to as Probably Approximately Correct (PAC) Learning. Conceptually, a class of (indicator) functions is PAC learnable if for any finitely many randomly generated examples there exists a suitable generalization function within the class. Blumer, Ehrenfeucht, Haussler and Warmuth [5] established in 1989 that an indicator function class is PAC learnable if and only if its VC dimension is finite. Their result is commonly referred to as the Fundamental Theorem of Statistical Learning. Both the argument in [5] and all following proofs of this theorem in the literature rely on “certain broadly applicable measurability conditions” (Ackerman, Asilis, Di, Freer and Tristan [1, page 1]). The VC dimension – introduced by Vapnik and Chervonenkis [^1] Due to different transliteration from the Cyrillic, various spellings of this name are common in Latin script. [31] in 1968 – is a parameter for measuring the combinatorial size of an indicator function class. Laying the basis for the Fundamental Theorem of Statistical Learning, Vapnik and Chervonenkis [32] proved in 1971 that finiteness of the VC dimension is a necessary and sufficient condition for uniform convergence of empirical means. In 1992, Laskowski [16] revealed a striking analogy between NIP formulas and classes having finite VC dimension. Here, NIP stands for ‘not the independence property’ (for details see Poizat [22, § 12.4]). This triangular relationship of the notions NIP, VC dimension and PAC learnability inspired a fruitful exchange between Model Theory and Statistical Learning Theory, as “connections between the VC theory and NIP have developed extensively with important notions from VC-theory adapted to the model-theoretic setting and vice versa” (Chase and Freitag [8, page 323]). For instance, Steinhorn [29, page 27] points out that “a definable family of definable sets $\mathcal{C}$ in an o-minimal structure is PAC learnable”. For drawing such connections special care has to be taken, as the Fundamental Theorem of Statistical Learning applies only to classes satisfying certain measurability conditions. In [5], a precise analysis of the measurability requirements is provided. The authors introduce the notion of well-behaved classes [^2] This is a measure-theoretic property that is developed in [5, Appendix A1]. We introduce a similar property in Section 3., to which the Fundamental Theorem of Statistical Learning applies. Although in [5, page 935] this property is described as “benign”, it requires “the measurability of complex sets of samples” (Ben-David, Benedek and Mansour [3, page 13]). Thus, the well-behavedness assumption constitutes a “complicated measurability constraint” (Ben-David, Itai and Kushilevitz [4, page 248]).
@@
The underlying PAC learning model considered in [5, § 2] relies on deterministic labels. More precisely, the samples on which the learning is based are generated according to distributions on the instance space and are equipped with labels that are generated according to a target function. In contrast, following more recent literature such as Shalev-Shwartz and Ben-David [25, § 3.2.1], we work with a model of PAC learning that is agnostic in various respects. This model omits target functions and is instead based on working with an arbitrary set of joint distributions on the sample space. Thus, the labels are not deterministic but instead also generated randomly. The agnostic PAC model therefore provides a more general and realistic description of formal learning (cf. Vidyasagar [34, § 3.3.1]). In particular, it is able to deal with noisy and corrupted data (cf. von Luxburg and Schölkopf [17, page 654 f.]). For further details see Remark 2.2, cf. also [5, Appendix A.3].
@@
Several expositions in Statistical Learning acknowledge that measurability subtleties demand a certain amount of attention (see e.g. Haussler [13, page 190], Mendelson and Smola [19, page 5], Pestov [20, page 1141] and Vidyasagar [34, page 56]). However, the available literature so far lacks a proof of the Fundamental Theorem of Statistical Learning that rigorously takes into account the measurability requirements of agnostic PAC learning. In this article, we fill this gap by presenting a self-contained proof that addresses the involved measurability issues in a detailed and precise manner.
@@
Our paper is structured as follows. In Section 2 we describe our learning framework and gather basic preliminaries from Statistical Learning Theory. In particular, we introduce the notions that are fundamental for the results we establish in Section 3: (agnostic) PAC Learning, Uniform Convergence Property and VC Dimension. Special emphasis is put on measure-theoretic aspects of these concepts. In Section 3 we present a sound statement as well as a detailed and self-contained proof of the Fundamental Theorem of Statistical Learning (Theorem 3.3). Its result relies on some measurability assumptions that are described in Definition 3.2. The proof comprises the verification of several implications (see Theorem 3.8, Theorem 3.9, Proposition 3.10 and Theorem 3.11). The technical heart is to prove that any hypothesis space with finite VC dimension has the uniform convergence property, provided certain measurability requirements are met (see Theorem 3.6 and Theorem 3.8). Section 4 initiates an examination of applications of the Fundamental Theorem of Statistical Learning in Model Theory. We introduce definable hypothesis spaces over first-order structures (Definition 4.1), which always have finite VC dimension if the underlying structure has NIP (Proposition 4.5). Our main theorem (Theorem 4.7) establishes sufficient conditions for the PAC learnability of hypothesis spaces defined over o-minimal expansions of the reals. We complete this work by gathering open questions in Section 5.
@@
Acknoledgements: We thank Matthias C. Caro, Salma Kuhlmann andTobias Sutter for several helpful discussions and for valuable advice on relevant literature. The second author would also like to extend special thanks to Salma Kuhlmann for the supervision of her doctoral project, which this work is part of. Additionally, the second author is grateful to David Pollard, Markus Kunze and Tim Seitz for pointing out various challenges related to measurability, clarifying some subtleties, and suggesting potential solutions during personal exchanges.
@@ 2 Framework and Preliminaries
    We assume that $\mathcal{P}_{\operatorname{fin}}(\mathcal{Z})\subseteq\Sigma_{\mathcal{Z}}$, or equivalently $\{z\}\in\Sigma_{\mathcal{Z}}$ for any $z\in\mathcal{Z}$. This is true if $\mathcal{Z}$ is a Hausdorff topological space – particularly, if $\mathcal{Z}$ is a metric space – and $\Sigma_{\mathcal{Z}}$ is the corresponding Borel $\sigma$–algebra [^4] Given a topological space $(\mathcal{Z},\tau)$, the corresponding Borel $\sigma$–algebra is the smallest $\sigma$–algebra containing $\tau$ as a subset.. Note further that our assumption implies $\Sigma_{\mathcal{Z}}=\mathcal{P}(\mathcal{Z})$ if $\mathcal{X}$ is countable. In the literature it is common to assume that $\mathcal{Z}$ is a metric space and that $\Sigma_{\mathcal{Z}}$ is the according Borel $\sigma$–algebra (see Haussler [14, page 91]).
@@
    The distributions in $\mathcal{D}$ represent the possible “states of nature” that might underlie the generation of the samples on which the learning is based (Haussler [14, page 93]). We can always choose to work with the set $\mathcal{D}^{\ast}$ containing all distributions to obtain distribution-free results.
@@
As we consider joint distributions $\mathbb{D}$ on the sample space $\mathcal{Z}$, it is possibly the case that for an instance $x\in\mathcal{X}$ both $(x,0)$ and $(x,1)$ have positive probability. In other words, the labels of instances are not deterministic or provided by a specific target function but are random themselves. This also allows us to deal with “noisy labels or corrupted data”. The underlying learning paradigm is sometimes referred to as agnostic (see [25, § 3.2.1]) or model-free (see [34, § 3.3], respectively). In contrast, Blumer, Ehrenfeucht, Haussler and Warmuth [5, § 2] work with a learning model that only involves distributions on the instance space $\mathcal{X}$. Thus, the instances $x_{1},\dots,x_{m}\in\mathcal{X}$ are drawn independently at random according to some probability distribution $\mathbb{P}$ on a $\sigma$–algebra on $\mathcal{X}$. The labels $y_{1},\dots,y_{m}$ are then generated according to a fixed target function $t\in\{0,1\}^{\mathcal{X}}$.
@@
$\mathbb{E}_{\underline{z}\sim\mathbb{D}^{m}}[\hat{\operatorname{er}}_{\underline{z}}(h)]=\frac{1}{m}\sum\limits_{i=1}^{m}\mathbb{E}_{z\sim\mathbb{D}}[\ell(h,z)]=\mathbb{E}_{z\sim\mathbb{D}}[\ell(h,z)]=\operatorname{er}_{\mathbb{D}}(h).$
Following [25, Definition 4.3] and [34, § 3.1.2], we now introduce the uniform convergence property, which guarantees that sample errors are “good” estimates for true errors.
@@
The VC dimension of $\mathcal{H}$, denoted $\mathrm{vc}(\mathcal{H})$, is the maximal size $d\in\mathbb{N}$ of a set $A\subseteq\mathcal{X}$ that can be shattered by $\mathcal{H}$. If the hypothesis space $\mathcal{H}$ can shatter sets of arbitrarily large size, then we say that $\mathcal{H}$ has infinite VC dimension and write $\mathrm{vc}(\mathcal{H})=\infty$.
@@
for $m\in\mathbb{N}$. The following result allows us to polynomially bound the values of the growth function using the VC dimension. It was established independently by Sauer [24, Theorem 1], Shelah [27], Vapnik and Chervonenkis [32, § 2], with motivation in different areas. As Shelah [27, page 254] also mentions that Perles was involved in the proof, it is sometimes called the Sauer–Shelah–Perles Lemma. Others refer to the result simply as Sauer’s Lemma.
@@
We finally specify the formal model of learning that we work with: Probably Approximately Correct (PAC) Learning. This concept, first introduced by Valiant [30], was modified and extended over time to also capture the agnostic setup that we deal with. The definition we use is a refined version of [2, Definition 2.1], [25, Definition 3.3] and [34, Definition 3.5], further inspired by [5, § 2].
@@
Note that the set $\{\underline{z}\in\mathcal{Z}^{m}\mid\operatorname{er}_{\mathbb{D}}(\mathcal{A}(\underline{z}))-\operatorname{opt}_{\mathbb{D}}(\mathcal{H})\leq\varepsilon\}$ is not necessarily contained in $\Sigma_{\mathcal{Z}}^{m}$ (see Example A.12). Further, we point out that in our model, PAC learnability of hypothesis spaces can be expressed with respect to any set $\mathcal{D}$ of distributions defined on the measurable space $(\mathcal{Z},\Sigma_{\mathcal{Z}})$. In the literature, however, PAC learnability is often defined in terms of the class of all (suitable) distributions (see e.g. [2, Definition 2.1] [25, Definition 3.3]).
@@
for any $m\in\mathbb{N}$ and $\underline{z}\in\mathcal{Z}^{m}$. Inspired by [34, § 3.3.2], we consider learning functions fulfilling a slightly weaker condition.
@@ 3 Fundamental Theorem of Statistical Learning
At this point, we emphasize that in the literature there are several extended versions of this theorem. These entail further equivalences and quantitative aspects such as sample complexity bounds (see e.g. [2, Theorem 5.5], [5, Theorem 2.1], [25, Theorem 6.7 and Theorem 6.8]). The main aim for the remainder of this section is to establish the Fundamental Theorem of Statistical Learning as stated above by virtue of the following more extended version:
@@
is a random variable [^7] The map $Y_{i}$ is a translation of the map $\theta_{i}\colon\mathcal{Z}^{m}\to\{0,1\},\ \underline{z}\mapsto\ell(h,z_{i})$, which is $\Sigma_{\mathcal{Z}}^{m}$–measurable, as $\Gamma(h)\in\Sigma_{\mathcal{Z}}$ implies $\theta_{i}^{-1}(\{0\})=\mathcal{Z}\times\dots\times\mathcal{Z}\times\underbrace{\Gamma(h)}_{i\text{--th position}}\times\mathcal{Z}\times\dots\times\mathcal{Z}\in\Sigma_{\mathcal{Z}}^{m}.$ with $\mathbb{E}_{\underline{z}\sim\mathbb{D}^{m}}[Y_{i}(\underline{z})]=0$ for any $i\in[m]$. Straightforward computations show that the random variables $Y_{1},\dots,Y_{m}$ are independent. Thus, we can apply Hoeffding’s Inequality (cf. Pollard [23, Appendix B, Corollary 3]) to obtain
@@
We next prove a result that shows how the growth function $\pi_{\mathcal{H}}$ can be used to bound the probability involved in the definition of the uniform convergence property. The proof requires the hypothesis space to be well-behaved, since its core lies in bounding the expected values of the random variables $U$ and $V$. The original proof of Vapnik and Chervonenkis is also based on the assumption that these maps are measurable (see [32, pages 265 and 268]). The proof presented here is based on the one of [25, Theorem 6.11]. We present here a modified and adapted version of [25, proof of Theorem 6.11], highlighting explicitly where the required measurability assumptions come into play.
@@
We have to show that $\mathbb{D}^{m}(\{\underline{z}\in\mathcal{Z}^{m}\mid U(\underline{z})\leq\varepsilon_{0}(m,\delta)\})\geq 1-\delta$. Since the random variable $U$ is non-negative, by Markov’s Inequality (see [25, Appendix B.1]) it suffices to establish that
@@
and it can be written as an average of independent random variables, each of which takes values in $[-1,1]$ (see Lemma A.1). Thus, we can apply Hoeffding’s Inequality (cf. Pollard [23, Appendix B, Corollary 3]) to obtain
@@
The arguments for proving the following theorem are inspired by the steps in [25, page 51].
@@
Note that this probability is well-defined, as $m\geq m_{0}\geq m_{\mathcal{H}}$ implies the measurability of the map $U(\mathcal{H},m,\mathbb{D})$ from Definition 2.4. Since $m\geq m_{0}^{(3)}$, applying [25, Lemma A.2] yields
@@
Blumer, Ehrenfeucht, Haussler and Warmuth [5] were the first to show that, given a hypothesis space $\mathcal{H}$ with finite VC dimension, any learning function $\mathcal{A}$ for $\mathcal{H}$, for which the determined hypothesis $\mathcal{A}(\underline{z})$ is consistent with the underlying multi-sample $\underline{z}$, is PAC. In our framework the consistency of $\mathcal{A}(\underline{z})$ with $\underline{z}$ can be translated into $\hat{\operatorname{er}}_{\underline{z}}(\mathcal{A}(\underline{z}))=0$. However, given $\underline{z}\in\mathcal{Z}^{m}$, in our agnostic setting there might not exist a hypothesis $h\in\mathcal{H}$ with $\hat{\operatorname{er}}_{\underline{z}}(h)=0$. We recall that to this end we introduced in Definition 2.8 the notion of learning functions that are nearly minimizing the sample error (NMSE). Adopting the computations in Vidyasagar [34, Theorem 3.2], we now show that, if a hypothesis space $\mathcal{H}$ has the uniform convergence property, then any such learning function is PAC.
@@
To complete the proof of the Fundamental Theorem of Statistical Learning, we now justify that hypothesis spaces with infinite VC dimension are not PAC learnable. Our arguments are guided by the proof of the No Free Lunch Theorem presented in [25, Theorem 5.1].
@@
Thus, applying Markov’s Inequality ([25, Appendix B.1]) yields

### v2

# Measurability in the Fundamental Theorem of Statistical Learning

@@ Abstract
The Fundamental Theorem of Statistical Learning states that a hypothesis space is PAC learnable if and only if its VC dimension is finite. For the agnostic model of PAC learning, the literature so far presents proofs of this theorem that often tacitly impose several measurability assumptions on the involved sets and functions. We scrutinize these proofs from a measure-theoretic perspective in order to explicitly extract the assumptions needed for a rigorous argument. This leads to a sound statement as well as a detailed and self-contained proof of the Fundamental Theorem of Statistical Learning in the agnostic setting, showcasing the minimal measurability requirements needed. As the Fundamental Theorem of Statistical Learning underpins a wide range of further theoretical developments, our results are of foundational importance: A careful analysis of measurability aspects is essential, especially when the theorem is used in settings where measure-theoretic subtleties play a role. We particularly discuss applications in Model Theory, considering NIP and o-minimal structures. Our main theorem presents sufficient conditions for the PAC learnability of hypothesis spaces defined over o-minimal expansions of the reals. This class of hypothesis spaces covers all artificial neural networks for binary classification that use commonly employed activation functions like ReLU and the sigmoid function.
Math Subject Classification (2020): Primary 68T05, 03C64; Secondary 28A05, 28A20, 03C98, 68T27, 12J15. Key Words: PAC Learning, VC Dimension, Uniform Convergence, Measurability, Borel Sets, NIP, o-Minimality. Corresponding Author: Laura Wirth, laura@uni-konstanz.de.
@@ 1 Introduction
In 1984, Valiant [32] introduced a formal model of learning based on randomly generated examples, which is referred to as Probably Approximately Correct (PAC) Learning. Conceptually, a class of (indicator) functions is PAC learnable if for any finitely many randomly generated examples there exists a suitable generalization function within the class. Blumer, Ehrenfeucht, Haussler and Warmuth [5] established in 1989 that an indicator function class is PAC learnable if and only if its VC dimension is finite. Their result is commonly referred to as the Fundamental Theorem of Statistical Learning. Both the argument in [5] and all following proofs of this theorem in the literature rely on “certain broadly applicable measurability conditions” (Ackerman, Asilis, Di, Freer and Tristan [1, page 1]). The VC dimension – introduced by Vapnik and Chervonenkis [^1] Due to different transliteration from the Cyrillic, various spellings of this name are common in Latin script. [33] in 1968 – is a parameter for measuring the combinatorial size of an indicator function class. Laying the basis for the Fundamental Theorem of Statistical Learning, Vapnik and Chervonenkis [34] proved in 1971 that finiteness of the VC dimension is a necessary and sufficient condition for uniform convergence of empirical means. In 1992, Laskowski [18] revealed a striking analogy between NIP formulas and classes having finite VC dimension. Here, NIP stands for ‘not the independence property’ (for details see Poizat [24, § 12.4]). This triangular relationship of the notions NIP, VC dimension and PAC learnability inspired a fruitful exchange between Model Theory and Statistical Learning Theory, as “connections between the VC theory and NIP have developed extensively with important notions from VC-theory adapted to the model-theoretic setting and vice versa” (Chase and Freitag [8, page 323]). For instance, Steinhorn [31, page 27] points out that “a definable family of definable sets $\mathcal{C}$ in an o-minimal structure is PAC learnable”. For drawing such connections special care has to be taken, as the Fundamental Theorem of Statistical Learning applies only to classes satisfying certain measurability conditions. In [5], a precise analysis of the measurability requirements is provided. The authors introduce the notion of well-behaved classes [^2] This is a measure-theoretic property that is developed in [5, Appendix A1]. We introduce a similar property in Section 3., to which the Fundamental Theorem of Statistical Learning applies. Although in [5, page 935] this property is described as “benign”, it requires “the measurability of complex sets of samples” (Ben-David, Benedek and Mansour [3, page 13]). Thus, the well-behavedness assumption constitutes a “complicated measurability constraint” (Ben-David, Itai and Kushilevitz [4, page 248]).
@@
The underlying PAC learning model considered in [5, § 2] relies on deterministic labels. More precisely, the samples on which the learning is based are generated according to distributions on the instance space and are equipped with labels that are generated according to a target function. In contrast, following more recent literature such as Shalev-Shwartz and Ben-David [27, § 3.2.1], we work with a model of PAC learning that is agnostic in various respects. This model omits target functions and is instead based on working with an arbitrary set of joint distributions on the sample space. Thus, the labels are not deterministic but instead also generated randomly. The agnostic PAC model therefore provides a more general and realistic description of formal learning (cf. Vidyasagar [36, § 3.3.1]). In particular, it is able to deal with noisy and corrupted data (cf. von Luxburg and Schölkopf [19, page 654 f.]). For further details see Remark 2.2, cf. also [5, Appendix A.3].
@@
Several expositions in Statistical Learning acknowledge that measurability subtleties demand a certain amount of attention (see e.g. Haussler [14, page 190], Mendelson and Smola [21, page 5], Pestov [22, page 1141] and Vidyasagar [36, page 56]). However, the available literature so far lacks a proof of the Fundamental Theorem of Statistical Learning that rigorously takes into account the measurability requirements of agnostic PAC learning. In this article, we fill this gap by presenting a self-contained proof that addresses the involved measurability issues in a detailed and precise manner.
@@
Our paper is structured as follows. In Section 2 we describe our learning framework and gather basic preliminaries from Statistical Learning Theory. In particular, we introduce the notions that are fundamental for the results we establish in Section 3: (agnostic) PAC Learning, Uniform Convergence Property and VC Dimension. Special emphasis is put on measure-theoretic aspects of these concepts. In Section 3 we present a sound statement as well as a detailed and self-contained proof of the Fundamental Theorem of Statistical Learning (Theorem 3.3). Its result relies on some measurability assumptions that are described in Definition 3.2. The proof comprises the verification of several implications (see Theorem 3.8, Theorem 3.9, Proposition 3.10 and Theorem 3.11). The technical heart is to prove that any hypothesis space with finite VC dimension has the uniform convergence property, provided certain measurability requirements are met (see Theorem 3.6 and Theorem 3.8). Section 4 initiates an examination of applications of the Fundamental Theorem of Statistical Learning in Model Theory. We introduce definable hypothesis spaces over first-order structures (Definition 4.1), which always have finite VC dimension if the underlying structure has NIP (Proposition 4.5). Our main theorem (Theorem 4.7) establishes sufficient conditions for the PAC learnability of hypothesis spaces defined over o-minimal expansions of the reals. This class of hypothesis spaces encompasses artificial neural networks whose activation functions are definable over the real exponential field (Example 4.2).We complete this work by gathering open questions in Section 5.
CRediT Authorship Contribution Statement: Lothar SebastianKrapp: conceptualization (supporting); funding acquisition (lead); investigation (supporting); project administration (lead); supervision (lead); writing – original draft (supporting), writing – review & editing (equal). Laura Wirth: conceptualization (lead); investigation (lead); writing – original draft (lead), writing – review & editing (equal).
@@
Acknowledgements: We thank Matthias C. Caro, Salma Kuhlmann andTobias Sutter for several helpful discussions and for valuable advice on relevant literature. The second author would also like to extend special thanks to Salma Kuhlmann for the supervision of her doctoral project, which this work is part of. Additionally, the second author is grateful to David Pollard, Itay Kaplan, Markus Kunze, Tim Seitz and Patrick Michalski for pointing out various challenges related to measurability, clarifying some subtleties, and suggesting potential solutions during personal exchanges.
Conflict of Interest: The authors declare no conflict of interest. The funders had no role in the design and conduct of the study; preparation, review, or approval of the manuscript; and decision to submit the manuscript for publication.
@@ 2 Framework and Preliminaries
    We assume that $\mathcal{P}_{\operatorname{fin}}(\mathcal{Z})\subseteq\Sigma_{\mathcal{Z}}$, or equivalently $\{z\}\in\Sigma_{\mathcal{Z}}$ for any $z\in\mathcal{Z}$. This is true if $\mathcal{Z}$ is a Hausdorff topological space – particularly, if $\mathcal{Z}$ is a metric space – and $\Sigma_{\mathcal{Z}}$ is the corresponding Borel $\sigma$–algebra [^4] Given a topological space $(\mathcal{Z},\tau)$, the corresponding Borel $\sigma$–algebra is the smallest $\sigma$–algebra containing $\tau$ as a subset.. Note further that our assumption implies $\Sigma_{\mathcal{Z}}=\mathcal{P}(\mathcal{Z})$ if $\mathcal{X}$ is countable. In the literature it is common to assume that $\mathcal{Z}$ is a metric space and that $\Sigma_{\mathcal{Z}}$ is the according Borel $\sigma$–algebra (see Haussler [15, page 91]).
@@
    The distributions in $\mathcal{D}$ represent the possible “states of nature” that might underlie the generation of the samples on which the learning is based (Haussler [15, page 93]). We can always choose to work with the set $\mathcal{D}^{\ast}$ containing all distributions to obtain distribution-free results.
@@
As we consider joint distributions $\mathbb{D}$ on the sample space $\mathcal{Z}$, it is possibly the case that for an instance $x\in\mathcal{X}$ both $(x,0)$ and $(x,1)$ have positive probability. In other words, the labels of instances are not deterministic or provided by a specific target function but are random themselves. This also allows us to deal with “noisy labels or corrupted data”. The underlying learning paradigm is sometimes referred to as agnostic (see [27, § 3.2.1]) or model-free (see [36, § 3.3], respectively). In contrast, Blumer, Ehrenfeucht, Haussler and Warmuth [5, § 2] work with a learning model that only involves distributions on the instance space $\mathcal{X}$. Thus, the instances $x_{1},\dots,x_{m}\in\mathcal{X}$ are drawn independently at random according to some probability distribution $\mathbb{P}$ on a $\sigma$–algebra on $\mathcal{X}$. The labels $y_{1},\dots,y_{m}$ are then generated according to a fixed target function $t\in\{0,1\}^{\mathcal{X}}$.
@@
$\mathbb{E}_{\underline{z}\sim\mathbb{D}^{m}}[\hat{\operatorname{er}}_{\underline{z}}(h)]=\frac{1}{m}\sum\limits_{i=1}^{m}\mathbb{E}_{z\sim\mathbb{D}}[\ell(h,z)]=\mathbb{E}_{z\sim\mathbb{D}}[\ell(h,z)]=\operatorname{er}_{\mathbb{D}}(h).\vspace{-0.7em}$
Following [27, Definition 4.3] and [36, § 3.1.2], we now introduce the uniform convergence property, which guarantees that sample errors are “good” estimates for true errors.
@@
The VC dimension of $\mathcal{H}$, denoted $\mathrm{vc}(\mathcal{H})$, is the maximal $d\in\mathbb{N}$ such that there exists a set $A\subseteq\mathcal{X}$ of size $d$ that is shattered by $\mathcal{H}$. If the hypothesis space $\mathcal{H}$ can shatter sets of arbitrarily large size, then we say that $\mathcal{H}$ has infinite VC dimension and write $\mathrm{vc}(\mathcal{H})=\infty$.
@@
for $m\in\mathbb{N}$. The following result allows us to polynomially bound the values of the growth function using the VC dimension. It was established independently by Sauer [26, Theorem 1], Shelah [29], Vapnik and Chervonenkis [34, § 2], with motivation in different areas. As Shelah [29, page 254] also mentions that Perles was involved in the proof, it is sometimes called the Sauer–Shelah–Perles Lemma. Others refer to the result simply as Sauer’s Lemma.
@@
We finally specify the formal model of learning that we work with: Probably Approximately Correct (PAC) Learning. This concept, first introduced by Valiant [32], was modified and extended over time to also capture the agnostic setup that we deal with. The definition we use is a refined version of [2, Definition 2.1], [27, Definition 3.3] and [36, Definition 3.5], further inspired by [5, § 2].
@@
Note that the set $\{\underline{z}\in\mathcal{Z}^{m}\mid\operatorname{er}_{\mathbb{D}}(\mathcal{A}(\underline{z}))-\operatorname{opt}_{\mathbb{D}}(\mathcal{H})\leq\varepsilon\}$ is not necessarily contained in $\Sigma_{\mathcal{Z}}^{m}$ (see Example A.12). [^7] For this reason, our notion of PAC learnability employs the computation of the inner measure, which can be computed also for sets that are not measurable. Further, we point out that in our model, PAC learnability of hypothesis spaces can be expressed with respect to any set $\mathcal{D}$ of distributions defined on the measurable space $(\mathcal{Z},\Sigma_{\mathcal{Z}})$. In the literature, however, PAC learnability is often defined in terms of the class of all (suitable) distributions (see e.g. [2, Definition 2.1] [27, Definition 3.3]).
@@
for any $m\in\mathbb{N}$ and $\underline{z}\in\mathcal{Z}^{m}$. Inspired by [36, § 3.3.2], we consider learning functions fulfilling a slightly weaker condition.
@@ 3 Fundamental Theorem of Statistical Learning
At this point, we emphasize that in the literature there are several extended versions of this theorem. These entail further equivalences and quantitative aspects such as sample complexity bounds (see e.g. [2, Theorem 5.5], [5, Theorem 2.1], [27, Theorem 6.7 and Theorem 6.8]). The main aim for the remainder of this section is to establish the Fundamental Theorem of Statistical Learning as stated above by virtue of the following more extended version:
@@
is a random variable [^8] The map $Y_{i}$ is a translation of the map $\theta_{i}\colon\mathcal{Z}^{m}\to\{0,1\},\ \underline{z}\mapsto\ell(h,z_{i})$, which is $\Sigma_{\mathcal{Z}}^{m}$–measurable, as $\Gamma(h)\in\Sigma_{\mathcal{Z}}$ implies $\theta_{i}^{-1}(\{0\})=\mathcal{Z}\times\dots\times\mathcal{Z}\times\underbrace{\Gamma(h)}_{i\text{--th position}}\times\mathcal{Z}\times\dots\times\mathcal{Z}\in\Sigma_{\mathcal{Z}}^{m}.$ with $\mathbb{E}_{\underline{z}\sim\mathbb{D}^{m}}[Y_{i}(\underline{z})]=0$ for any $i\in[m]$. Straightforward computations show that the random variables $Y_{1},\dots,Y_{m}$ are independent. Thus, we can apply Hoeffding’s Inequality (cf. Pollard [25, Appendix B, Corollary 3]) to obtain
@@
We next prove a result that shows how the growth function $\pi_{\mathcal{H}}$ can be used to bound the probability involved in the definition of the uniform convergence property. The proof requires the hypothesis space to be well-behaved, since its core lies in bounding the expected values of the random variables $U$ and $V$. The original proof of Vapnik and Chervonenkis is also based on the assumption that these maps are measurable (see [34, pages 265 and 268]). The proof presented here is based on the one of [27, Theorem 6.11]. We present here a modified and adapted version of [27, proof of Theorem 6.11], highlighting explicitly where the required measurability assumptions come into play.
@@
We have to show that $\mathbb{D}^{m}(\{\underline{z}\in\mathcal{Z}^{m}\mid U(\underline{z})\leq\varepsilon_{0}(m,\delta)\})\geq 1-\delta$. Since the random variable $U$ is non-negative, by Markov’s Inequality (see [27, Appendix B.1]) it suffices to establish that
@@
and it can be written as an average of independent random variables, each of which takes values in $[-1,1]$ (see Lemma A.1). Thus, we can apply Hoeffding’s Inequality (cf. Pollard [25, Appendix B, Corollary 3]) to obtain
@@
The arguments for proving the following theorem are inspired by the steps in [27, page 51].
@@
Note that this probability is well-defined, as $m\geq m_{0}\geq m_{\mathcal{H}}$ implies the measurability of the map $U(\mathcal{H},m,\mathbb{D})$ from Definition 2.4. Since $m\geq m_{0}^{(3)}$, applying [27, Lemma A.2] yields
@@
Blumer, Ehrenfeucht, Haussler and Warmuth [5] were the first to show that, given a hypothesis space $\mathcal{H}$ with finite VC dimension, any learning function $\mathcal{A}$ for $\mathcal{H}$, for which the determined hypothesis $\mathcal{A}(\underline{z})$ is consistent with the underlying multi-sample $\underline{z}$, is PAC. In our framework the consistency of $\mathcal{A}(\underline{z})$ with $\underline{z}$ can be translated into $\hat{\operatorname{er}}_{\underline{z}}(\mathcal{A}(\underline{z}))=0$. However, given $\underline{z}\in\mathcal{Z}^{m}$, in our agnostic setting there might not exist a hypothesis $h\in\mathcal{H}$ with $\hat{\operatorname{er}}_{\underline{z}}(h)=0$. We recall that to this end we introduced in Definition 2.8 the notion of learning functions that are nearly minimizing the sample error (NMSE). Adopting the computations in Vidyasagar [36, Theorem 3.2], we now show that, if a hypothesis space $\mathcal{H}$ has the uniform convergence property, then any such learning function is PAC.
@@
To complete the proof of the Fundamental Theorem of Statistical Learning, we now justify that hypothesis spaces with infinite VC dimension are not PAC learnable. Our arguments are guided by the proof of the No Free Lunch Theorem presented in [27, Theorem 5.1].
@@
Thus, applying Markov’s Inequality ([27, Appendix B.1]) yields